# A two-phase model of galaxy formation: III. The formation of globular clusters

### v1

@@ 2.3 Star formation, SMBH growth and feedback
where $\alpha_{\rm agn,f}=10^{-3}$, $\beta_{\rm agn,s}=1.8$, and $M_{\rm bh}$ and $M_{\rm g}$ are the masses of central SMBH and the gas within the galaxy (equations describing their evolution are given below), respectively. Modeled in this way, the effects of the SN and AGN feedback are constrained by the gravitational potential of the host halo, which is needed to reproduce the observed $M_{*}$-$M_{\rm v}$ relation at the low-mass ($\approx 10^{8}{\rm M_{\odot}}$) end (Yang et al., 2012a) and the scaling relation of $M_{\rm bh}$-$M_{\rm*,bulge}$ (Graham & Sahu, 2023). The fraction of ejected gas is set to be $f_{\rm ej}=0.75$ for the fast phase. This choice was made in Paper-I to ensure that the gas fraction in a feedback-regulated galaxy is comparable to the typical spin, $\lambda\approx 0.04$, so that a switch from bulge formation to disk formation can occur after the fast phase. For the slow phase, we set $f_{\rm ej}=1$. Note that in the slow phase, the hot ISM component represented by $\Delta M_{\rm g,prev}$ is degenerate with the remaining part of $\Delta M_{\rm g,sf}$ after the consumption of star formation, so that $f_{\rm ej}$ is degenerate with $F_{\rm sn}$ and $F_{\rm agn}$, and the star formation efficiency of the star-forming gas, $\Delta M_{\rm g,sf}$. Once the star formation efficiency is fixed, the choice of $f_{\rm ej}$ does affect the gas fraction in the galaxy. The roughly constant gas fraction shown by the green curve in the bottom left panel of Fig. 2 indicates that gas fraction predicted by our choice for $f_{\rm ej}$ is reasonable.
@@
Galaxy mergers are modeled by adding a fraction, $f_{\rm merge}$, of the stellar mass of the satellite galaxy into the central galaxy. The remaining stars are assumed to be tidally stripped or scattered into the halo to become intra-cluster light (ICL). The fiducial value, $f_{\rm merge}=0.1$ is taken from Model-II of L14. However, we note that this value is highly uncertain because of the complicated details of mergers and the degeneracy with other model parameters. Given the uncertainty, whenever necessary, we will bracket $M_{*}$ by showing two extremes, one only includes stars that have formed in the main branch ($f_{\rm merge}=0$), and the other include stars that have formed in both the main branch and all the side branches ($f_{\rm merge}=1$).
@@ 2.4 Metal enrichment
Gas-phase metallicity, $Z\equiv M_{\rm Z}/M_{\rm g}$, of galaxies is modeled according to the ‘gas-regulator’ scenario (Lilly et al., 2013; Peng et al., 2015; Wang & Lilly, 2021), as our modeling of the SFR is based on the amount of available gas and fits in the framework of the gas regulator. However, certain modifications are needed to account for uncertainties in the metal enrichment related to the formation of GCs. One uncertainty arises from the possibility that high-$z$ galaxies may be fed by cold streams, as suggested by numerical simulations (e.g. Keres et al., 2005; Ceverino et al., 2010; Danovich et al., 2015; Mandelker et al., 2018; Lapiner et al., 2023). Cold streams with small cross section can transport pristine gas into the galaxy, leading to the formation of metal-poor stars, as suggested by observations (Lin et al., 2023), analytical estimations (Li et al., 2023) and hydrodynamical simulations (Mandelker et al., 2018). Another uncertainty comes from the expectation that high-$z$ galaxies are much more clumpy and that their star formation is more bursty than in the local Universe, as detailed in Paper-I. Metal yields from stars at high $z$ may thus be coupled to the ISM and IGM in a way very different from that in a more steady and quieter environment expected at low $z$. One way to deal with these uncertainties is to directly use the observed relations of the metallicity with other galaxy properties that can be modeled reliably. This approach, however, is limited by the size and quality of observational data, the selection bias, and uncertainties in, e.g., tracers and calibrations used to infer the metallicity. It is also not guaranteed that the observed relation can be applied to modeling the GCs, as the environment of GCs may differ significantly from the average over an entire galaxy. Indeed, the GC catalog produced by post-processing the FIRE-2 zoom-in simulation, which is claimed to be able to reproduce the observed stellar mass-metallicity relation up to $z=12$ (Marszewski et al., 2024), misses the old (born at $z>3$), metal-poor population observed in the Milky Way (Grudić et al., 2023).
@@
where the three terms on the right-hand side are referred to as the inflow, yield and locked metal masses, respectively. Thus, $Z_{\rm IGM}$ is the metallicity of the inflow gas, $y_{\rm eff}$ is the effective oxygen yield that incorporates the intrinsic yield of stars, the escape of decoupled fraction, the ejection by stellar feedback, and the recycling due to cooling; and $R$ is the fraction of mass returned to the ISM by stellar evolution. Note that gas ejection and recycling has been absorbed into the effective yield, $y_{\rm eff}$, so that this factor accounts for the net effects of ejection and recycling. Each of the terms in the above equation has uncertainties. To proceed, we make the following assumptions:
@@
    A constant returning fraction, $R=0.4$, is adopted by assuming a Chabrier IMF (Bruzual & Charlot, 2003). Instantaneous recycling is also assumed here, because the time scale of stellar evolution is much shorter than the dynamical time scale of the halo – roughly $80\%$ of the oxygen is released during the first $10\,{\rm Myr}$ by the Type-II SNe (Maiolino & Mannucci, 2019).
@@
$\log\frac{Z}{Z_{\odot}}=0.6\log(\frac{M_{*}}{10^{9}{\rm M_{\odot}}})-0.55\log(1+z)-0.5\,,$ (27)
@@
The model of galaxy formation described in this section focuses on the global properties of galaxies. As described earlier, this provides environmental conditions for small-scale processes related to the formation of GCs. A consequence of this hierarchical strategy is that the spatial distribution of GCs resembles those of the halo and galaxies at large scales, with modifications due to refined processes operating on small scales. In the next section, we will introduce the model of GC formation under the constraints provided by the model of galaxies described above.
@@ 3 The formation of globular clusters
A main conclusion of our two-phase model (Paper-I) is the formation of sub-structures, referred to as sub-clouds, once the SGC becomes self-gravitating. In the fast phase, the SGC is expected to be gas-rich and turbulent due to the fast accretion of halo. High-speed winds from SN and AGN can also contribute to the turbulence (Ma et al., 2020; Mercedes-Feliz et al., 2023). Significant compression of the sub-clouds is expected if the turbulence is supersonic, which may produce dense sub-clouds for the formation of GCs we are interested in here. Despite the complexity of the details, it is generally believed that the properties of sub-clouds in a turbulent ISM follow ‘quasi-universal’ laws and distributions, as suggested by both observations (Larson, 1981; Lada & Dame, 2020) and numerical simulations (Burkhart et al., 2017; Ma et al., 2020; Grudić et al., 2021; Appel et al., 2022; Buck et al., 2022; Kiihne et al., 2023). In this section, we sample the sub-cloud population using these universal distributions under the constraints given by the global properties of SGCs. We then derive the condition for star formation in each sub-clouds, taking into account both external and internal feedback effects on the sub-clouds. We demonstrate that the condition for a sub-cloud to form a GC leads to two distinct channels for GC formation. Finally, we model the disruption and survival of GCs, which is needed to make predictions for the observed population.
@@ 3.1 The formation of sub-clouds
The maximum mass of sub-clouds at formation is set by the Jeans mass. In the limit of atomic cooling, the temperature $T_{\rm cool}\approx 10^{4}{\rm K}$, corresponding to a sound speed of $c_{\rm s}\approx 10\,{\rm km/s}$, and the Jeans mass is
@@
where $n_{\rm gas}=10^{3.5}\,{\rm cm}^{-3}\,n_{\rm 3.5}=1\,{\rm cm}^{-3}\,n_{\rm 0.0}$, $v_{\rm eff}=(c_{\rm s}^{2}+v_{\rm turb}^{2})^{1/2}=10\,{\rm{km\,s^{-1}}}\,v_{\rm eff,10}$, and $v_{\rm turb}$ is the turbulence velocity. The precise mass distribution of sub-clouds depends on the details of how SGCs fragment, and therefore on the temperature, turbulence velocity, and density of individual SGCs.
The mass function of sub-clouds can be inferred by the mass function of newly born massive star clusters, which follows a power-law distribution with an index $\approx-2$, as found by observations in the local Universe (Fall & Chandar, 2012; Adamo et al., 2020) and the cosmic noon (Dessauges-Zavadsky & Adamo, 2018). This is also confirmed by high-resolution simulations of individual GMCs (Grudić et al., 2021), and is usually interpreted as the result of the scale-free fragmentation of a turbulence-dominated ISM (Larson, 1981). In this paper, we adopt the following initial mass function for sub-clouds:
@@
with $\beta_{\rm sc,m}=-2$. The truncation mass $M_{\rm sc,t}=10^{6.5}{\rm M_{\odot}}$ is set to reflect the Jeans mass given above and the mass limit on the most massive star clusters found in observations (e.g. Krumholz et al., 2019, see also Fig. 10). The lower and upper bounds of $M_{\rm sc}$ are set to be $M_{\rm sc,min}=10^{4}{\rm M_{\odot}}$ and $M_{\rm sc,max}=10^{8}{\rm M_{\odot}}$, respectively. Convergence tests against these choices are made in §4.2.
The presence of supersonic shocks (with Mach number $\mathcal{M}$ > 1) can raise the density of sub-clouds and significantly change the density distribution of sub-clouds. This, coupled with non-linear star formation and feedback processes, can modify the mass spectrum. Major sources of the turbulence may be summarized as follows:
@@
    Global shocks. These arise during the formation of an SGC when the inflow gas reaches the core of the host halo, interacts with itself and collides with the existing gas. This is especially important for halos with fast accretion and effective cooling, where cold streams penetrate the halo (corresponding to the Q1 phase described in Paper-I, ), as suggested by numerical simulations (e.g. Ceverino et al., 2010; Danovich et al., 2015; Mandelker et al., 2018). A global shock can also arise when angular momentum, turbulence and thermal motion of the SGC can no longer support its gravity and it collapses (e.g. Dekel et al., 2009; Dekel & Burkert, 2014; Zolotov et al., 2015; Tacchella et al., 2016; Latif et al., 2022; Ji & Giavalisco, 2023; Dekel et al., 2023). This is expected to be an important source of turbulence for a gas-rich, slow-rotating SGC (corresponding to the Q2 phase). Gas-rich mergers also produce global shocks and generate gravitational instability due to the pile-up of gas and loss of angular momentum. Such a global shock can propagate outward across the entire SGC, compress the gas, and produce high-density sub-structures.
@@
    Local interactions. These occur when a sub-cloud grows via absorbing surrounding gas, or mergers with nearby sub-clouds. The release of gravitational energy can be effectively dissipated by atomic cooling, so that this process proceeds rapidly and is roughly isothermal. The process continues until sub-clouds become too small and dense enough for them to collide frequently, or until the driving force of the global turbulence disappears. Numerical simulations have shown that such local interactions are critical for the formation of massive and dense sub-clouds (e.g. Latif et al., 2022; Ma et al., 2020).

### v2

@@ 2.3 Star formation, SMBH growth and feedback
where $\alpha_{\rm agn,f}=10^{-3}$, $\beta_{\rm agn,s}=1.8$, and $M_{\rm bh}$ and $M_{\rm g}$ are the masses of central SMBH and the gas within the galaxy (equations describing their evolution are given below), respectively. Modeled in this way, the effects of the SN and AGN feedback are constrained by the gravitational potential of the host halo, which is needed to reproduce the observed $M_{*}$-$M_{\rm v}$ relation at the low-mass ($\approx 10^{8}{\rm M_{\odot}}$) end (Yang et al., 2012a) and the scaling relation of $M_{\rm bh}$-$M_{\rm*,bulge}$ (Graham & Sahu, 2023). The fraction of ejected gas is set to be $f_{\rm ej}=0.75$ for the fast phase. This choice was made in Paper-I to ensure that the gas fraction in a feedback-regulated galaxy is comparable to the typical spin, $\lambda\approx 0.04$, so that a switch from bulge formation to disk formation can occur after the fast phase. For the slow phase, we set $f_{\rm ej}=1$. Note that in the slow phase, the hot ISM component represented by $\Delta M_{\rm g,prev}$ is degenerate with the remaining part of $\Delta M_{\rm g,sf}$ after the consumption of star formation, so that $f_{\rm ej}$ is degenerate with $F_{\rm sn}$ and $F_{\rm agn}$, and the star formation efficiency of the star-forming gas, $\Delta M_{\rm g,sf}$. Once the star formation efficiency is fixed, the choice of $f_{\rm ej}$ does affect the gas fraction in the galaxy. The roughly constant gas fraction shown by the green curve in the bottom left panel of Fig. 2 indicates a gas fraction of $M_{\rm g}/M_{\rm v}\approx\lambda\approx 0.04$. Such a gas fraction allows a transition from bulge to disk at the end of the fast phase, and the formation of a stable disk in the slow phase (see §3.2 of Mo et al., 1998), suggesting that the gas fraction predicted by our choice for $f_{\rm ej}$ is reasonable.
@@
Galaxy mergers are modeled by adding a fraction, $f_{\rm merge}$, of the stellar mass of the satellite galaxy into the central galaxy. The remaining stars are assumed to be tidally stripped or scattered into the halo to become intracluster light (ICL). The fiducial value, $f_{\rm merge}=0.1$ is taken from Model-II of L14. However, we note that this value is highly uncertain because of the complicated details of mergers and the degeneracy with other model parameters. This value also depends on the definition of the galaxy stellar mass in observations, i.e. the boundary used to separate galactic stars from intergalactic stars and the aperture used in deriving the stellar mass. To quantify this uncertainty, we bracket $M_{*}$ by two extremes: one only includes stars that have formed in the main branch ($f_{\rm merge}=0$), and the other includes stars that have formed in both the main branch and all the side branches ($f_{\rm merge}=1$). The resulted $M_{*}$-$M_{\rm v}$ relations for these two extremes, shown by the green curves in Fig. 12 (ses §4.4 for details), suggest a significant contribution by ex-situ stars at $M_{\rm v}\gtrsim 10^{13}{\rm M_{\odot}}$. Interpretations of our results involving $M_{*}$ should thus take into account this uncertainty. When counting the total number and mass of GCs in a galaxy (see §4.1), we instead include all GCs formed in satellites that have merged into the central, based on the fact that observations of GCs usually adopt a large aperture (e.g. Dornan & Harris, 2023).
@@ 2.4 Metal enrichment
Gas-phase metallicity, $Z\equiv M_{\rm Z}/M_{\rm g}$, of galaxies is modeled according to the ‘gas-regulator’ scenario (Lilly et al., 2013; Peng et al., 2015; Wang & Lilly, 2021), as our modeling of the SFR is based on the amount of available gas and fits in the framework of the gas regulator. However, certain modifications are needed to account for uncertainties in the metal enrichment related to the formation of GCs. One uncertainty arises from the possibility that high-$z$ galaxies may be fed by cold streams, as suggested by numerical simulations (e.g. Keres et al., 2005; Ceverino et al., 2010; Danovich et al., 2015; Mandelker et al., 2018; Lapiner et al., 2023). Cold streams with small cross section can transport pristine gas into the galaxy, leading to the formation of metal-poor stars, as suggested by observations (Lin et al., 2023), analytical estimations (Li et al., 2024) and hydrodynamical simulations (Mandelker et al., 2018). Another uncertainty comes from the expectation that high-$z$ galaxies are much more clumpy and that their star formation is more bursty than in the local Universe, as detailed in Paper-I. Metal yields from stars at high $z$ may thus be coupled to the ISM and IGM in a way very different from that in a more steady and quieter environment expected at low $z$. One way to deal with these uncertainties is to directly use the observed relations of the metallicity with other galaxy properties that can be modeled reliably. This approach, however, is limited by the size and quality of observational data, the selection bias, and uncertainties in, e.g., tracers and calibrations used to infer the metallicity. It is also not guaranteed that the observed relation can be applied to modeling the GCs, as the environment of GCs may differ significantly from the average over an entire galaxy. Indeed, the GC catalog produced by post-processing the FIRE-2 zoom-in simulation, which is claimed to be able to reproduce the observed stellar mass-metallicity relation up to $z=12$ (Marszewski et al., 2024), misses the old (born at $z>3$), metal-poor population observed in the Milky Way (Grudić et al., 2023).
@@
where the three terms on the right-hand side are referred to as the inflow, yield and locked metal masses, respectively. Thus, $Z_{\rm IGM}$ is the metallicity of the inflow gas, $y_{\rm eff}$ is the effective oxygen yield that incorporates the intrinsic yield of stars, the escape of decoupled fraction, the ejection by stellar feedback, and the recycling due to cooling; Note that gas ejection and recycling has been absorbed into the effective yield, $y_{\rm eff}$, so that this factor accounts for the net effects of ejection and recycling. Each of the terms in the above equation has uncertainties. To proceed, we make the following assumptions:
@@
    A constant returning fraction $R$ (Eq. 17) is adopted. Instantaneous recycling is also assumed here, because the timescale of stellar evolution is much shorter than the dynamical timescale of the halo – roughly $80\%$ of the oxygen is released during the first $10\,{\rm Myr}$ by the Type-II SNe (Maiolino & Mannucci, 2019).
@@
$\log\frac{Z}{\,{\rm Z}_{\odot}}=0.6\log(\frac{M_{*}}{10^{9}{\rm M_{\odot}}})-0.55\log(1+z)-0.5\,,$ (27)
@@
The model of galaxy formation described in this section focuses on the global properties of galaxies. As described earlier, this provides environmental conditions for small-scale processes related to the formation of GCs. A consequence of this hierarchical strategy is that the spatial distribution of GCs resembles those of the halo and galaxies at large scales, with modifications due to refined processes operating on small scales. In the next section, we will introduce the model of GC formation under the constraints provided by the model of galaxies described above. Note that this top-down approach does not include the details of matter and energy recycling from stars to the gas environment, and thus misses the diversity in the feedback produced by different stellar populations. For example, the IMF in low-metallicity environments is inferred to be top-heavy (Stacy et al., 2016; Latif et al., 2022b; Klessen & Glover, 2023), which can alter the evolution paths of individual stars and thus lead to different metal yields (Portinari et al., 1998). More exotic feedback mechanisms can also be present, such as kilonovae (Rosswog et al., 2013; Metzger, 2017) and pair-instability supernovae (Woosley et al., 2002; Janka, 2012). The inclusion of these populations and related processes may be important in the early Universe, and further exploration is needed in this area.
@@ 3 The formation of globular clusters
A main conclusion of our two-phase model (Paper-I) is the formation of sub-structures, referred to as sub-clouds, once the SGC becomes self-gravitating. In the fast phase, the SGC is expected to be gas-rich and turbulent due to the fast accretion of halo. High-speed winds from SN and AGN can also contribute to the turbulence (Ma et al., 2020; Mercedes-Feliz et al., 2023). Significant compression of the sub-clouds is expected if the turbulence is supersonic, which may produce dense sub-clouds for the formation of GCs we are interested in here. Despite the complexity of the details, it is generally believed that the properties of sub-clouds in a turbulent ISM follow ‘quasi-universal’ laws and distributions, as suggested by both observations (Larson, 1981; Lada & Dame, 2020) and numerical simulations (Burkhart et al., 2017; Ma et al., 2020; Grudić et al., 2021; Appel et al., 2022; Buck et al., 2022; Kiihne et al., 2023). In this section, we sample the sub-cloud population using these universal distributions under the constraints given by the global properties of SGCs. We then derive the condition for star formation in each sub-clouds, taking into account both external and internal feedback effects on the sub-clouds. We demonstrate that the condition for a sub-cloud to form a GC leads to two distinct channels for GC formation. Finally, we model the disruption and survival of GCs, which is needed to make predictions for comparisons with observations.
@@ 3.1 The formation of sub-clouds
At formation, an SGC is shock-heated by the release of gravitational potential energy to a temperature determined by the virial temperature of the host halo. Gas cooling in the SGC is efficient, as expected from the high gas density and the high rate of atomic cooling expected at the relevant temperature. The rapid drop of the gas temperature reduces the Jeans mass, and can trigger fragmentation of the SGC into sub-structures. The decline of the temperature is expected to be stalled at $T_{\rm cool}\approx 10^{4}\,{\rm K}$ where the cooling rate starts to drop precipitously (e.g. Smith et al., 2008; Wiersma et al., 2009) and the cooling timescale becomes long. Thermal motion of gas particles in a cloud fragment can thus operate over a sufficient period of time to suppress density fluctuations in the fragment, stabilizing it and preventing it from further fragmentation. The Jeans mass at $T_{\rm cool}$ therefore sets a characteristic mass for the fragmentation of SGCs. In this paper, we refer to these self-bound fragments, whose initial temperature is $T_{\rm cool}$, as sub-clouds. Each of these sub-clouds is expected to feature a thermally-supported internal structure and a mass spectrum truncated by the Jeans mass in the high-mass end. The characteristic temperature, $T_{\rm cool}\approx 10^{4}{\rm K}$, corresponds to a sound speed of $c_{\rm s}\approx 10\,{\rm km/s}$, and the Jeans mass is
@@
where $n_{\rm gas}=10^{3.5}\,{\rm cm}^{-3}\,n_{\rm 3.5}=1\,{\rm cm}^{-3}\,n_{\rm 0.0}$, $v_{\rm eff}=(c_{\rm s}^{2}+v_{\rm turb}^{2})^{1/2}=10\,{\rm{km\,s^{-1}}}\,v_{\rm eff,10}$, and $v_{\rm turb}$ is the turbulence velocity. By definition, $v_{\rm turb}<c_{\rm s}$, so that $v_{\rm eff}\sim c_{\rm s}$. We note that the meaning of the sub-clouds so defined should be distinguished from that of the GMCs in the literature, with the latter usually referring to molecular gas clouds of diverse properties.
The precise mass distribution of sub-clouds depends on the details of how SGCs fragment, and therefore on the temperature, turbulence velocity and density of individual SGCs, as well as on the magnetic and tidal fields they reside in. We do not attempt to trace these details here, but adopt an empirical approach instead. We assume that the mass function of sub-clouds below the truncation mass follows a power-law distribution with an index $\beta_{\rm sc,m}$, as expected from scale-free fragmentation of a turbulence-dominated ISM (Larson, 1981). The value of $\beta_{\rm sc,m}$ is inferred to be $\approx-2$, as suggested by the index of the mass function of newly-born massive star clusters in the local Universe (Zhang & Fall, 1999; Fall & Chandar, 2012; Adamo et al., 2020) and in the cosmic noon (Dessauges-Zavadsky & Adamo, 2018). The star formation efficiency of massive and dense sub-clouds hosting massive and dense star clusters is expected to be close to unity (see §3.4), and thus the mass spectrum of sub-clouds is expected to be similar to the mass function of star clusters. This value of $\beta_{\rm sc,m}$ is also reproduced by high-resolution simulations of individual GMCs (Grudić et al., 2021). With the consideration above, we model the initial mass function for sub-clouds as
@@
with $\beta_{\rm sc,m}=-2$. The truncation mass $M_{\rm sc,t}=10^{6.5}{\rm M_{\odot}}$ is set to reflect the Jeans mass given above and the mass limit on the most massive star clusters found in observations (e.g. Mok et al. 2019; Krumholz et al. 2019; see also Fig. 11). The lower and upper bounds of $M_{\rm sc}$ are set to be $M_{\rm sc,min}=10^{4}{\rm M_{\odot}}$ and $M_{\rm sc,max}=10^{8}{\rm M_{\odot}}$, respectively. Convergence tests against these choices are made in §4.2.
The presence of supersonic shocks (with Mach number $\mathcal{M}$ > 1) can raise the density of sub-clouds and significantly change the density distribution of sub-clouds. As explained in §3.3 of Paper-I, the compressive nature of supersonic turbulence is critical to make the sub-clouds dense enough to survive the drag force by ram pressure and collisions with other sub-clouds. A lack of supersonic turbulence, such as the case where a dynamically-cold disk forms and steady star formation ensues, is expected to terminate the generation of sub-clouds at $T_{\rm cool}$ before gas cooling at lower temperature sets a new mass scale, $M_{\rm sc,t}$, that is lower than the original truncation mass. Sub-cloud formation in this regime of weak-turbulence may be related to local instabilities, such as those induced by spiral arms and SN feedback, and/or to external perturbations, such as those from interactions with nearby galaxies. On the other hand, sub-clouds in strong turbulence are clustered (Li et al., 2020). They may thus coalesce to form more massive clouds, raising the value of $M_{\rm sc,t}$, or form massive star clusters if rapid star formation can consume the gas before the coalescence (see e.g. Fig. 9 of Ma et al., 2020). Such a variation of $M_{\rm sc,t}$ is evident from observations (see §2.2 of Krumholz et al., 2019, for a review), and may be modeled by following the evolution of the turbulence in the SGC. Alternatively, one may start from a parametric model that relates $M_{\rm sc,t}$ to the galactic environment, such as the gas density, composition and star formation rate density, with necessary calibration from hydrodynamic simulations and observations. Lacking such a model, we adopt a constant value of $M_{\rm sc,t}$ for simplicity.
The density distribution of sub-clouds in supersonic turbulence, coupled with non-linear star formation and feedback processes, can modify the mass spectrum of sub-clouds and thus that of star clusters formed within them. Major sources of the turbulence may be summarized as follows:
@@
    Global shocks. These arise during the formation of an SGC when the inflow gas reaches the core of the host halo, interacts with itself and collides with the existing gas. This is especially important for halos with fast accretion and effective cooling, where cold streams penetrate the halo (corresponding to the Q1 phase described in Paper-I, ), as suggested by numerical simulations (e.g. Ceverino et al., 2010; Danovich et al., 2015; Mandelker et al., 2018). A global shock can also arise when angular momentum, turbulence and thermal motion of the SGC can no longer support its gravity and it collapses (e.g. Dekel et al., 2009; Dekel & Burkert, 2014; Zolotov et al., 2015; Tacchella et al., 2016; Latif et al., 2022a; Ji & Giavalisco, 2023; Dekel et al., 2023). This is expected to be an important source of turbulence for a gas-rich, slow-rotating SGC (corresponding to the Q2 phase). Gas-rich mergers also produce global shocks and generate gravitational instability due to the pile-up of gas and loss of angular momentum. Such a global shock can propagate outward across the entire SGC, compress the gas, and produce high-density sub-structures.
@@
    Local interactions. These occur when a sub-cloud grows via absorbing surrounding gas, or mergers with nearby sub-clouds. The release of gravitational energy can be effectively dissipated by the effective cooling above $10^{4}\,{\rm K}$, so that this process proceeds rapidly and is roughly isothermal. The process continues until sub-clouds become too small and dense enough for them to collide frequently, or until the driving force of the global turbulence disappears. Numerical simulations have shown that such local interactions are critical for the formation of massive and dense sub-clouds (e.g. Latif et al., 2022a; Ma et al., 2020).